An undiscovered planet, many light-years from Earth, has one moon which has a nearly circular periodic orbit. If the distance from the center of the moon to the surface of the planet is 2.00000 × 105 km and the planet has a radius of 3225 km and a mass of 6.85 × 1022 kg, how long (in days) does it take the moon to make one revolution around the planet. The gravitational constant is 6.67 × 10-11 N·m2/kg2.
Answer #1

d = 2x10^5 km , R = 3225 km , M =6.85x10^22 kg, G =6.67x10^-11 N.m^2/kg^2

a =d +R = (2x105)=(0.03225x105) =2.03225x105km

a =2.03225x108m

From kepler's third law

T2 = (4\pi2/GM)a3

T2 =[ 4(3.14)2 (2.03225x108)3]/[(6.67x10-11)(6.85x1022)]

T2 =72.6x1012

T = 8.5x106s

1day = 24x60x60 s

1s = 1.1574x10-5 days

T = 8.5x106(1.1574x10-5 days)

T = 98.38days